OpenStudy (anonymous):

You start out with $20 and then spend money in a store where every item is $3. Y = -3x + 20. Type how to graph the equation using the y-intercept and slope. Be sure to tell where you place the first point and how to find the location of the second point.

OpenStudy (texaschic101):

y = -3x + 20 In y = mx + b form, the slope is in the m position, and the y intercept is in the b position. So our y int = (0,20) Now plot your point on (0,20) and since your slope is -3,or -3/1, we go down 3 spaces (because it is negative) and to the right 1 space,(plot your point) and down 3, and to the right 1,(plot your point).....keep doing this. y = -3x + 20 --- find your x axis by subbing in 0 for y 0 = -3x + 20 3x = 20 x = 20/3 (for graphing purposes, it is 6 2/3) By finding the x axis (20/3,0), this tells you where the line your were working on will cross the x line

OpenStudy (anonymous):

Thank you so much